OpenStudy (shubhamsrg):

infinite summation : 1 + (1/3) + (1.3 / 3.6) + (1.3.5 / 3.6.9) + (1.3.5.7 / 3.6.9.12) ........ = ?

OpenStudy (shubhamsrg):

so far, my progress has been this much : i started my series from (1/3) ... denom. of nth term is easily (3^n)*(n!) num. of nth term will be 1.3.5.7............ =(1.2.3.4.5...........)/(2.4.6.8........) =( (2n)! )/(2^n * n! ) thus nth term will be (2n)! / (n!)^2 * 6^n or C(2n,n) / 6^n we have to fine its summation from n = 1 to infinity.. but am clueless after this.. will this be even helpful ?

OpenStudy (experimentx):

that's central binomial coefficient ... google for something like it.

OpenStudy (shubhamsrg):

well i have googled over,,i got this solution,,but couldnt understand as to why we did that.. it said to compare the summation with (1+x)^n and solve ultimately,,and to my surprise, i got the ans.. if anyone can explain why ?

OpenStudy (experimentx):

it's better to use the generating function directly http://mathworld.wolfram.com/images/equations/CentralBinomialCoefficient/NumberedEquation1.gif

OpenStudy (shubhamsrg):

didnt get you ?

OpenStudy (experimentx):

http://mathworld.wolfram.com/CentralBinomialCoefficient.html put x=1/6 and get it's value

OpenStudy (shubhamsrg):

ohh,,nice !! where does this come from ?

OpenStudy (experimentx):